# Practical security of twin-field quantum key distribution under wavelength-switching attack

### v1

@@ Abstract
The twin-field class quantum key distribution (TF-class QKD) has experimentally demonstrated the ability to surpass the fundamental rate-distance limit without requiring a quantum repeater, as a revolutional milestone. In TF-class QKD implementation, an optical phase-locked loop (OPLL) structure is commonly employed to generate a reference light with correlated phase, ensuring coherence of optical fields between Alice and Bob. In this configuration, the reference light, typically located in the untrusted station Charlie, solely provides wavelength reference for OPLL and does not participate in quantum-state encoding. However, the reference light may open a door for Eve to enter the source stations that are supposed to be well protected. Here, by identifying vulnerabilities in the OPLL scheme, we propose and demonstrate a wavelength-switching attack on a TF-class QKD system. This attack involves Eve deliberately manipulating the wavelength of the reference light to increase mean photon number of prepared quantum states, while maintaining stable interference between Alice and Bob as required by TF-class QKD protocols. The maximum observed increase in mean photon number is 8.7%, which has been theoretically proven to compromise the security of a TF-class QKD system. Moreover, we have shown that with well calibration of the modulators, the attack can be eliminated. Through this study, we highlight the importance of system calibration in the practical security in TF-class QKD implementation.
I Introduction
@@
Fortunately, a recent protocol called twin-field QKD (TF QKD) has been proposed, surpassing the fundamental rate-distance limit and revolutionizing the scalability of both key rate and channel loss to be $O(\sqrt{\eta})$ [6]. TF QKD leverages the advantage that two distant peers, Alice and Bob, transmit two optical fields with correlated phase to conduct single-photon interference at an intermediate station named Charlie. This implies that only one photon from either Alice or Bob needs to arrive and be detected at Charlie in order to generate a secret key, while the other peer transmits vacuum states that are inherently immune to channel loss. In addition to overcoming transmission limits, TF QKD also inherits the advantages of measurement-device-independent (MDI) QKD [7, 8], where Charlie can be fully untrusted as an intermediate station. The security of TF QKD against general attacks has been theoretically proven, establishing a solid foundation for its implementation [9, 10, 11, 12, 13, 14]. Following this theoretical development, experimental demonstrations have shown that TF-class QKD outperforms the PLOB bound [15, 16, 17, 18, 19, 20, 21, 22], achieving record-breaking distances up to $1002~{}\mathrm{k}\mathrm{m}$ [22, 23]. These achievements highlight the feasibility of implementing intercity QKD without relying on trusted relays.
@@
However, the practical security performance, another crucial aspect of QKD, has not yet been investigated in the context of this novel TF-class QKD. Regarding its implementation, it is quite challenging to generate identical optical fields with correlated phase by two separate parties as required by the protocol. This unique requirement introduces a distinct implementation scheme that may also introduce unforeseen security vulnerabilities. In initial proof-of-principle experiments [6, 24], the light source was placed at an untrusted location and transmitted to Alice and Bob. Unfortunately, this approach is evidently insecure since Eve can manipulate the quantum signal encoding key information by controlling the light source. To mitigate this vulnerability and strictly adhere to the structure of TF QKD protocol, subsequent experiments equipped each Alice and Bob with separate lasers. However, this mitigation strategy gives rise to a new challenge - ensuring that Alice’s and Bob’s lasers produce optical fields with identical phase correlation. To address this concern, achieving wavelength consistency of the two separate laser sources is accomplished through the laser injection technique. In this technique, either Charlie applies a seed laser to independently inject into Alice’s and Bob’s laser sources [25, 18], or Alice employs an ultra-stable light source to directly transmit the seed light beam to Bob [15, 17]. However, this scheme potentially introduces a vulnerability for Eve to gain access to the assumed-to-be-protected source station [26, 27] - Eve could manipulate the seed light via Charlie or exploit the public channel for similar laser-seeding attacks [28, 29, 30]. Moreover, TF QKD schemes might also be susceptible to Trojan-horse attacks. Therefore, ensuring practical security in TF-class QKD systems is more intricate compared to previous QKD implementations due to their unique realization schemes and techniques that require heightened attention.
@@
In addition to the aforementioned known threats on these schemes, an advanced technique known as optical phase-locked loop (OPLL) has been extensively employed in recent implementations to achieve persistent stable interference by generating identical optical fields between Alice and Bob [15, 17, 18, 31, 20, 21, 22, 32]. In this approach, Alice’s and Bob’s lasers are synchronized with a reference laser source for both frequency and phase [18, 20, 21, 22]. While this method enhances the stability of correlated phases in two optical fields and supports the realization of a secure interference TF-class QKD implementation, its practical security aspects remain unexplored.
@@
The structure of this paper is as follows. In Sec. II, we introduce the idea of TF-class QKD and discuss the implementation of OPLL in the TF-class QKD system. Building upon this, we propose a wavelength-switching attack specifically designed for this system. In Sec. III, we perform characterization tests on the acousto-optic modulators (AOM) device used in the OPLL structure, followed by an experimental demonstration of the wavelength-switching attack on the TF-class QKD system. Section IV analyzes the security of TF-class QKD systems based on the experimental results obtained from the wavelength-switching attack. In Sec. V, we discuss some potential countermeasures against the proposed hacking strategy. Finally, this study is summarized and concluded in Sec. VI.
II The idea of wavelength-switching attack
@@
Figure 1: Experimental schematic diagram and testing results. (a) Schematic diagram illustrating the structure of the optical phase-locked loop (OPLL) in a TF-class quantum key distribution (QKD) system, along with the corresponding wavelength-switching attack. The red and blue connecting lines represent optical fiber and electrical cables, respectively. EDFAs refer to erbium-doped fiber amplifiers, FLC stands for frequency-locking controller, BS denotes beam splitter, PMBS represents polarization-maintaining beam splitter, and PZT refers to piezoelectric transducer. (b) Interference measurements obtained by Charlie in the TF QKD system under both normal operation conditions and during the presence of a wavelength-switching attack at $f_{\Delta}=30~{}\mathrm{M}\mathrm{Hz}$. (c) Optical power measurements recorded at AOM2’s output while subjected to a wavelength-switching attack with $r_{s}=100~{}\mathrm{k}\mathrm{Hz}$.
II.1 TF-class QKD based on OPLL
@@
To achieve correlated phase between two optical fields, Alice and Bob employ the OPLL structure in TF-class QKD systems for wavelength and phase synchronization, as depicted in Fig. 1(a). The OPLL comprises a frequency-locking controller (FLC), piezoelectric transducer (PZT), AOM, and heterodyne frequency measurement. Charlie utilizes an ultra-stable laser to generate a reference light, which is subsequently transmitted to Alice and Bob. This reference light does not carry any encoding information but can be amplified through erbium-doped fiber amplifiers (EDFAs) during transmission. Upon receiving the reference light, Alice/Bob employs heterodyne detection to determine the frequency/phase difference between her/his laser and the reference light. Specifically, a photodetector and phase discriminator detect the frequency/phase deviation, which is then processed by a proportional integrator (PI) controller to adjust Alice’s/Bob’s laser frequency/phase to match that of the reference light. This adjustment is achieved by controlling PZTs and two inserted AOMs in their respective lasers. The PZTs are used for adjusting initial frequencies and constraining slow frequency drifts while AOMs compensate for fast frequency shifts. The coordinated operation of PZTs and AOMs enables precise frequency synchronization between Alice’s and Bob’s lasers in TF-class QKD systems.
II.2 The working principle of wavelength-switching attack
@@
In the TF-class QKD system, exemplified by Sec. II.1, the OPLL structure facilitates precise adjustment of Alice’s or Bob’s laser wavelength through coordinated actions of AOM and PZT. Our primary focus lies in elucidating the operational procedure of OPLL and exploiting a loophole in the AOM within this configuration to control the output intensities of quantum states prepared by Alice and Bob. Specifically, we observe that any sudden fluctuations in wavelength trigger prompt corrective shifts by the AOM. However, an unforeseen imperfection arises from the dependence of insertion loss on modulation frequency $f_{aom}$, which is further illustrated in Appendix A.
Based on the aforementioned concept, we propose a wavelength-switching attack for the TF-class QKD system, as depicted by the red area in Fig. 1(a). In this attack scenario, a wavelength-switching module consisting of an AOM and an arbitrary waveform generator (AWG) is inserted at the output of the reference laser. Consequently, the AOM induces frequency shifts between $f_{aom1}$ and $f_{aom2}$, while the AWG controls the rate of wavelength switching ($r_{s}$). By carefully adjusting these parameters, Eve ensures that only AOM operates without any adjustment on PZT in Alice’s/Bob’s OPLL module. Subsequently, she transmits a reference light with a pattern of switched wavelengths to both Alice and Bob. Upon receiving this reference light, Alice and Bob measure its heterodyne frequency relative to their local light source. The measurement reveals two peak values corresponding to different wavelengths of the reference light, prompting FLC to adjust AOM levels in Alice/Bob accordingly. In essence, Eve manipulates the wavelength of reference light to trigger wavelength switching in Alice’s/Bob’s local light sources. Importantly, this wavelength switching also reduces insertion loss in AOMs leading to increased intensity of quantum states prepared by Alice/Bob. If not monitored properly by Alice and Bob through their output light signals, this increased intensity could potentially enable Eve to eavesdrop on their secret key.
@@ III.1 The characteristics of the AOM
To investigate the characteristics of the AOM, a comprehensive test was conducted to examine its response under various input conditions, with a specific focus on two key dimensions: carrier frequency and driving voltage. Regarding the test on carrier frequency, a sinusoidal driving signal with frequencies $f_{aom}$ ranging from 00 to $240~{}\mathrm{M}\mathrm{Hz}$ was injected into the AOM. The initial laser wavelength, $\lambda_{in}=v/f_{\rm{in}}=1550~{}\mathrm{n}\mathrm{m}$ where $v$ represents the speed of light in optical fiber and $f_{i}n$ is the frequency. Upon receiving the driving signal, the AOM effectively modulates the light’s wavelength to $\lambda_{out}=v/\left(f_{\rm{in}}+f_{\rm{aom}}\right)$. By continuously varying $f_{aom}$ during the experiment, accurate characterization of how different values of $f_{aom}$ corresponded to changes in output power was obtained. During testing on driving voltage, a fixed-frequency sinusoidal driving signal at $200$ $\mathrm{M}\mathrm{Hz}$ was applied to the AOM while systematically adjusting peak-to-peak voltages within a range of $1$ $\mathrm{V}$ to $15$ $\mathrm{V}$. This meticulous adjustment allowed for establishing an empirical relationship between variations in input voltage levels and corresponding changes in output efficiency for precise performance evaluation.
The relationship between AOM attenuation and frequency modulation is illustrated in Fig. 2(a). In this subfigure, the black solid curve represents simulation results, while the red dots correspond to experimental measurements. The observed attenuation spans a range from over $50~{}\mathrm{d}\mathrm{B}$ to a few decibels as a function of modulation frequency. Notably, the minimum measured attenuation value for the AOM employed in this study is $3.57~{}\mathrm{d}\mathrm{B}$, occurring at a modulation frequency of $195~{}\mathrm{M}\mathrm{Hz}$. This variation in attenuation primarily stems from the effective coupling of light within the AOM device and its piezoelectric transducer efficiency. Additionally, imperfect coupling may contribute to discrepancies between experimental data (red dots) and simulation results (black curve). Disregarding coupling efficiency reveals that higher modulation frequencies result in increased ultrasonic power, directly influencing output efficiency. For comprehensive details on the working principle and model of AOM, please refer to  Appendix A.
@@
Thus, in the TF-class QKD system under wavelength-switching attack, the attenuation of an AOM is influenced by both frequency modulation and drive voltage. For instance, when $f_{aom1}=200~{}\mathrm{M}\mathrm{Hz}$, $f_{aom2}=180~{}\mathrm{M}\mathrm{Hz}$ and, thus, $f_{\Delta}=20~{}\mathrm{M}\mathrm{Hz}$, Fig. 2 (a) shows that frequency modulation results in an increase in attenuation by $0.69~{}\mathrm{d}\mathrm{B}$. Meanwhile, Fig. 2(b) illustrates that the driving voltage is increased from $8.95~{}\mathrm{V}$ (black dot) to $10.56~{}\mathrm{V}$ (red upward triangle, high voltage), resulting in a $1.02~{}\mathrm{d}\mathrm{B}$ decrease in attenuation. The combined effect of these two factors ultimately leads to an overall attenuation reduction of $0.33~{}\mathrm{d}\mathrm{B}$ for the AOM. This reduction in AOM attenuation corresponds to the increase in Alice’s output optical power as depicted in Fig. 1 (c). When the AOM2 (AOM3) work frequency returns to $200~{}\mathrm{M}\mathrm{Hz}$, Fig. 2(b) also illustrates that the driving voltage reduces to $8.68~{}\mathrm{V}$ (red downward triangle, low voltage), resulting in a $0.14~{}\mathrm{d}\mathrm{B}$ increase in attenuation. Demonstration of wavelength-switching attack with more details are presented in the following subsections.
III.2 Testing of wavelength-switching attack

### v2

@@ Abstract
The twin-field class quantum key distribution (TF-class QKD) has experimentally demonstrated the ability to surpass the fundamental rate-distance limit without requiring a quantum repeater, as a revolutional milestone. In TF-class QKD implementation, an optical phase-locked loop (OPLL) structure is commonly employed to generate a reference light with correlated phase, ensuring coherence of optical fields between Alice and Bob. In this configuration, the reference light, typically located in the untrusted station Charlie, solely provides wavelength reference for OPLL and does not participate in quantum-state encoding. However, the reference light may open a door for Eve to enter the source stations that are supposed to be well protected. Here, by identifying vulnerabilities of an acousto-optic modulator (AOM) in the OPLL scheme, we propose and demonstrate a wavelength-switching attack on a TF-class QKD system. This attack involves Eve deliberately manipulating the wavelength of the reference light to increase mean photon number of prepared quantum states, while maintaining stable interference between Alice and Bob as required by TF-class QKD protocols. The maximum observed increase in mean photon number is 8.7%, which has been theoretically proven to compromise the security of a TF-class QKD system. Moreover, we have shown that with well calibration of the modulators, the attack can be eliminated. Through this study, we highlight the importance of system calibration in the practical security in TF-class QKD implementation.
1 Introduction
@@
Fortunately, a recent protocol called twin-field QKD (TF QKD) has been proposed, surpassing the fundamental rate-distance limit and revolutionizing the scalability of both key rate and channel loss to be $O(\sqrt{\eta})$ [6]. TF QKD leverages the advantage that two distant peers, Alice and Bob, transmit two optical fields with correlated phase to conduct single-photon interference at an intermediate station named Charlie. This implies that only one photon from either Alice or Bob needs to arrive and be detected at Charlie in order to generate a secret key, while the other peer transmits vacuum states that are inherently immune to channel loss. In addition to overcoming transmission limits, TF QKD also inherits the advantages of measurement-device-independent (MDI) QKD [7, 8], where Charlie can be fully untrusted as an intermediate station. The security of TF QKD against general attacks has been theoretically proven, establishing a solid foundation for its implementation [9, 10, 11, 12, 13, 14]. Following this theoretical development, experimental demonstrations have shown that TF-class QKD outperforms the PLOB bound [15, 16, 17, 18, 19, 20, 21, 22, 23], achieving record-breaking distances up to $1002~{}\mathrm{k}\mathrm{m}$ [23, 24]. These achievements highlight the feasibility of implementing intercity QKD without relying on trusted relays.
@@
However, the practical security performance, another crucial aspect of QKD, has not yet been investigated in the context of this novel TF-class QKD. Regarding its implementation, it is quite challenging to generate identical optical fields with correlated phase by two separate parties as required by the protocol. This unique requirement introduces a distinct implementation scheme that may also introduce unforeseen security vulnerabilities. In initial proof-of-principle experiments [6, 25], the light source was placed at an untrusted location and transmitted to Alice and Bob. Unfortunately, this approach is evidently insecure since Eve can manipulate the quantum signal encoding key information by controlling the light source. To mitigate this vulnerability and strictly adhere to the structure of TF QKD protocol, subsequent experiments equipped each Alice and Bob with separate lasers. However, this mitigation strategy gives rise to a new challenge - ensuring that Alice’s and Bob’s lasers produce optical fields with identical phase correlation. To address this concern, achieving wavelength consistency of the two separate laser sources is accomplished through the laser injection technique. In this technique, Charlie applies a seed laser to independently inject into Alice’s and Bob’s laser sources [26, 20]. However, this scheme potentially introduces a vulnerability for Eve to gain access to the assumed-to-be-protected source station [27, 28] - Eve could manipulate the seed light via Charlie or exploit the public channel for similar laser-seeding attacks [29, 30, 31]. Moreover, TF QKD schemes might also be susceptible to Trojan-horse attacks and induced-photorefractive attacks [32, 33]. Therefore, ensuring practical security in TF-class QKD systems is more intricate compared to previous QKD implementations due to their unique realization schemes and techniques that require heightened attention.
@@
In addition to the aforementioned known threats on these schemes, an advanced technique known as optical phase-locked loop (OPLL) has been extensively employed in recent implementations to achieve persistent stable interference by generating identical optical fields between Alice and Bob [15, 34, 16, 17, 19, 20, 21, 22, 23, 35]. In this approach, Alice’s and Bob’s lasers are synchronized with a reference laser source for both frequency and phase [20, 21, 22, 23]. While this method enhances the stability of correlated phases in two optical fields and supports the realization of a secure interference TF-class QKD implementation, its practical security aspects remain unexplored.
@@
2 Results
2.1 The idea of wavelength-switching attack
@@
Figure 1: Experimental schematic diagram and testing results. (a) Schematic diagram illustrating the structure of the OPLL in a TF-class QKD system, along with the corresponding wavelength-switching attack. The red and blue connecting lines represent optical fiber and electrical cables, respectively. The frequency-locking controller (FLC) modulates the operation of Alice’s/Bob’s piezoelectric transducer (PZT) and acousto-optic modulator (AOM2/AOM3) to ensure the coherence of optical field between Alice and Bob. The optical intensity of Alice’s/Bob’s laser is measured by photodetector PD5/PD6, whose bandwidth is $500~{}\mathrm{M}\mathrm{Hz}$, connected to an oscilloscope with sampling rate of 80 GSa/s. EDFAs, erbium-doped fiber amplifiers; BS, beam splitter; Het, heterodyne measurement; PMBS, polarization-maintaining beam splitter. (b) Interference measurements obtained by Charlie in the TF QKD system under both normal operation conditions and during the presence of a wavelength-switching attack at $f_{\Delta}=30~{}\mathrm{M}\mathrm{Hz}$. (c) Optical power measurements recorded at AOM2’s output while subjected to a wavelength-switching attack with the rate of wavelength switching $r_{s}=100~{}\mathrm{k}\mathrm{Hz}$.
2.1.1 TF-class QKD based on OPLL
@@
To achieve correlated phase between two optical fields, Alice and Bob employ the OPLL structure in TF-class QKD systems for wavelength and phase synchronization, as depicted in Fig. 1(a). The OPLL comprises a frequency-locking controller (FLC), piezoelectric transducer (PZT), acousto-optic modulator (AOM), and heterodyne measurement. Charlie utilizes an ultra-stable laser to generate a reference light, which is subsequently transmitted to Alice and Bob. This reference light does not carry any encoding information but can be amplified through erbium-doped fiber amplifiers (EDFAs) during transmission. Upon receiving the reference light, Alice/Bob employs heterodyne measurement to determine the frequency/phase difference between her/his laser and the reference light. Specifically, a photodetector and phase discriminator detect the frequency/phase deviation, which is then processed by a proportional integrator (PI) controller to adjust Alice’s/Bob’s laser frequency/phase to match that of the reference light. This adjustment is achieved by controlling PZTs and two inserted AOMs in their respective lasers. The PZTs are used for adjusting initial frequencies and constraining slow frequency drifts while AOMs compensate for fast frequency shifts. When there is a change in the wavelength of the reference light, the FLC rapidly adjusts the frequency modulation via the AOM. Following this, to maintain the AOM in their optimal operating state, the FLC gradually alters the frequency modulation of the AOM back to its central value, while concurrently modifying the PZT to compensate for the frequency discrepancy. The coordinated operation of PZTs and AOMs enables precise frequency synchronization between Alice’s and Bob’s lasers in TF-class QKD systems.
2.1.2 The working principle of wavelength-switching attack
@@
In the TF-class QKD system, exemplified by Section 2.1.1, the OPLL structure facilitates precise adjustment of Alice’s or Bob’s laser wavelength through coordinated actions of AOM and PZT. Our primary focus lies in elucidating the operational procedure of OPLL and exploiting a loophole in the AOM within this configuration to control the output intensities of quantum states prepared by Alice and Bob. Specifically, we observe that any sudden fluctuations in wavelength trigger prompt corrective shifts by the AOM. However, an unforeseen imperfection arises from the dependence of insertion loss on modulation frequency $f_{aom}$, which is further illustrated in Section 4.1.
Based on the aforementioned concept, we propose a wavelength-switching attack for the TF-class QKD system, as depicted by the red area in Fig. 1(a). In this attack scenario, a wavelength-switching module consisting of an AOM1 and an arbitrary waveform generator (AWG) is inserted at the output of the reference laser, as shown in the red box in Fig. 1(a). Consequently, the AOM induces frequency shifts between $f_{aom1}$ and $f_{aom2}$, while the AWG controls the rate of wavelength switching ($r_{s}$). By carefully adjusting these parameters, Eve ensures that only AOM operates without any adjustment on PZT in Alice’s/Bob’s OPLL module. Subsequently, she transmits a reference light with a pattern of switched wavelengths to both Alice and Bob. Upon receiving this reference light, Alice and Bob measure its heterodyne frequency relative to their local light source. The measurement reveals two peak values corresponding to different wavelengths of the reference light, prompting FLC to adjust AOM levels in Alice/Bob accordingly. In essence, Eve manipulates the wavelength of reference light to trigger wavelength switching in Alice’s/Bob’s local light sources. Importantly, this wavelength switching also reduces insertion loss in AOMs leading to increased intensity of quantum states prepared by Alice/Bob. If not monitored properly by Alice and Bob through their output light signals, this increased intensity could potentially enable Eve to eavesdrop on their secret key.
@@ 2.2.1 The characteristics of the AOM
To investigate the characteristics of the AOM, a comprehensive test was conducted to examine its response under various input conditions, with a specific focus on two key dimensions: carrier frequency and driving voltage. Regarding the test on carrier frequency, a constant driving voltage of $15.84~{}\mathrm{V}$ was applied to the AOM, and a sinusoidal driving signal with frequencies faom ranging from 00 to $240~{}\mathrm{M}\mathrm{Hz}$ was injected into it. The initial laser wavelength, $\lambda_{\rm{in}}=v/f_{\rm{in}}=1550~{}\mathrm{n}\mathrm{m}$ where $v$ represents the speed of light in optical fiber and $f_{\rm{in}}$ is the frequency. Upon receiving the driving signal, the AOM effectively modulates the light’s wavelength to $\lambda_{out}=v/\left(f_{\rm{in}}+f_{\rm{aom}}\right)$. By continuously varying $f_{aom}$ during the experiment, accurate characterization of how different values of $f_{aom}$ corresponded to changes in output power was obtained. During testing on driving voltage, a fixed-frequency sinusoidal driving signal at $200$ $\mathrm{M}\mathrm{Hz}$ was applied to the AOM while systematically adjusting peak-to-peak voltages within a range of $1$ $\mathrm{V}$ to $15$ $\mathrm{V}$. This meticulous adjustment allowed for establishing an empirical relationship between variations in input voltage levels and corresponding changes in output efficiency for precise performance evaluation.
The relationship between AOM attenuation and frequency modulation is illustrated in Fig. 2(a). In this subfigure, the black solid curve represents simulation results, while the red dots correspond to experimental measurements. The observed attenuation spans a range from over $50~{}\mathrm{d}\mathrm{B}$ to a few decibels as a function of modulation frequency. Notably, the minimum measured attenuation value for the AOM employed in this study is $3.57~{}\mathrm{d}\mathrm{B}$, occurring at a modulation frequency of $195~{}\mathrm{M}\mathrm{Hz}$. This variation in attenuation primarily stems from the effective coupling of light within the AOM device and its piezoelectric transducer efficiency. Additionally, imperfect coupling may contribute to discrepancies between experimental data (red dots) and simulation results (black curve). Disregarding coupling efficiency reveals that higher modulation frequencies result in increased ultrasonic power, directly influencing output efficiency. For comprehensive details on the working principle and model of AOM, please refer to  Section 4.1.
@@
Thus, in the TF-class QKD system under wavelength-switching attack, the attenuation of an AOM is influenced by both frequency modulation and drive voltage. For instance, when $f_{aom1}=200~{}\mathrm{M}\mathrm{Hz}$, $f_{aom2}=180~{}\mathrm{M}\mathrm{Hz}$ and, thus, $f_{\Delta}=20~{}\mathrm{M}\mathrm{Hz}$, Fig. 2 (a) shows that frequency modulation results in an increase in attenuation by $0.69~{}\mathrm{d}\mathrm{B}$. Meanwhile, Fig. 2(b) illustrates that the driving voltage is increased from $8.95~{}\mathrm{V}$ (black dot) to $10.56~{}\mathrm{V}$ (red upward triangle, high voltage), resulting in a $1.02~{}\mathrm{d}\mathrm{B}$ decrease in attenuation. The combined effect of these two factors ultimately leads to an overall attenuation reduction of $0.33~{}\mathrm{d}\mathrm{B}$ for the AOM, which is consistent with the measured value, as indicated in Fig. 2(c) when $f_{\Delta}=20~{}\mathrm{M}\mathrm{Hz}$. When the AOM2 (AOM3) work frequency returns to $200~{}\mathrm{M}\mathrm{Hz}$, Fig. 2(b) also illustrates that the driving voltage reduces to $8.68~{}\mathrm{V}$ (red downward triangle, low voltage), resulting in a $0.14~{}\mathrm{d}\mathrm{B}$ increase in attenuation. Demonstration of wavelength-switching attack with more details are presented in the following subsections.
2.2.2 Testing of wavelength-switching attack